A P50,000 debt is to be repaid at the end of one year. The debtor establishes a sinking

fund that earns 8% interest compounded quarterly. Construct a sinking fund

schedule.


1
Expert's answer
2021-11-12T11:10:51-0500

Solution:

Amount to be repaid (FV)=50000

Interest rate on sinking fund = 8%

Per quarter interest rate (i)=APR/Number of compounding in year

=8%/4 = 2%

Debt repaid in 1 year

Number of quarters in 1 year (n)=1*4 = 4

Sinking fund amount

"= FV\times \dfrac{i}{(1+i)^n-1}

\\=50000\times \dfrac{2\%}{(1+2\%)^4-1}

\\=12131.18763"

Sinking fund schedule is created as below

Amount of each sinking fund =12131.18763

Interest each period = Beginning balance * quarterly interest rate

Ending balance = Beginning balance + interest + Sinking fund deposit

Sinking fund schedule: